# DSOcc: Leveraging Depth Awareness and Semantic Aid to Boost Camera-Based 3D Semantic Occupancy Prediction

### v1

@@ C.1 Supplementary Ablation Study
Image Segementation Impact. To assess the effect of image segmentation on overall performance, we replaced MSeg [25] with the MobileNet [19] and RegSeg [24] trained on the Cityscapes dataset [9]. As illustrated in Table 11, although these models exhibits lower accuracy and suboptimal class alignment compared to MSeg, it still raises the baseline in mIoU. This improvement confirms the effectiveness of semantic-aided method.
@@
Depth Estimation Error Impact. To assess how depth estimation error affect our model performance, we randomly added noise to the estimated depth to simulate real-world estimation errors. Using the performance of MobileStereoNet [46] on the KITTI 2015 validation set [15] (EPE = 0.79 px) as the baseline, we tested our method across various error ratios. As illustrated in Fig. 7, our method achieves mIoU scores of 16.06 % and 14.96 % when adding one and two times noise, respectively, demonstrating the robustness of soft occupancy confidence.
Frame Design and Efficiency. To evaluate our frame design and efficiency, especially in the final feature fusion, we introduce a comparison called ‘Three Fus’, which fuses lifted image, segmentation, and depth features via dual interaction. Table 12 reports the inference time of our voxel fusion module alongside other voxel-based learning modules. Our depth-aware fusion attains a higher mIoU than ThreeFus. Our voxel fusion processes dense voxels in 177.07 ms, comparable to the sparse-voxel modules used in VoxFormer [32] and SGN [42].
Soft Occupancy Confidence Calculation. To compare different methods for computing soft occupancy confidence, we evaluated the performance of tanh and sigmoid functions, each combined with a range-mapping operation to normalize inputs from , into . As illustrated in Table 13, these functions yield comparable results and effectively meet our objectives. However, our exponential-based function achieved a highest mIoU, indicating its superior performance.
| Ours | Tanh | Sigmoid | |
| mIoU | 18.02 | 14.85 | 14.99 |
| IoU | 44.41 | 42.31 | 39.57 |
@@
Depth Using. To assess the advantage of our depth embedding, we compare it against two alternatives: using depth truncation to represent explicit occupancy state and to learn features for occupied voxels [32, 53], and using depth to generate geometric cues that interact with image features [23, 56]. As shown in Table 14, our depth embedding reduces false-negative occupancy predictions and thereby improves mIoU from 13.35 to 15.00. By lifting depth and image features into a spatially aligned 3D representation, our method outperforms the depth-interaction approach, since aligned features enable more accurate fusion.
@@ C.2 Supplementary Qualitative Results
To demonstrate the robustness of DSOcc, we present additional visualization results in Fig. 8. DSOcc effectively identifies each vehicle in front of the ego vehicle. It can also detect small objects, such as pedestrians, cyclists, and poles. Furthermore, DSOcc provides robust scene completion when navigating intersections.
@@ Appendix D Limitation
As shown in Fig. 9, when the vehicle rounds a corner, our method sometimes fails to reconstruct the side view because of missing camera features and abrupt changes in vehicle calibration. Our method recognizes this region as the flat other-ground class. This issue could be mitigated by incorporating additional camera viewpoints or by integrating other modalities, such as LiDAR.

### v3

@@ C.1 Supplementary Ablation Study
Image Segementation Impact. To assess the effect of image segmentation on overall performance, we replaced MSeg [21] with the MobileNet [15] and RegSeg [20] trained on the Cityscapes dataset [8]. As illustrated in Table C1, although these models exhibits lower accuracy and suboptimal class alignment compared to MSeg, it still raises the baseline in mIoU. This improvement confirms the effectiveness of semantic-aided method.
@@
Depth Estimation Error Impact. To assess how depth estimation error affect our model performance, we randomly added noise to the estimated depth to simulate real-world estimation errors. Using the performance of MobileStereoNet [42] on the KITTI 2015 validation set [12] (EPE = 0.79 px) as the baseline, we tested our method across various error ratios. As illustrated in Fig. C1, our method achieves mIoU scores of 16.06 % and 14.96 % when adding one and two times noise, respectively, demonstrating the robustness of soft occupancy confidence.
@@
Depth Using. To assess the advantage of our depth embedding, we compare it against two alternatives: using depth truncation to represent explicit occupancy state and to learn features for occupied voxels [27, 47], and using depth to generate geometric cues that interact with image features [19, 50]. As shown in Table C2, our depth embedding reduces false-negative occupancy predictions and thereby improves mIoU from 13.35 to 15.00. By lifting depth and image features into a spatially aligned 3D representation, our method outperforms the depth-interaction approach, since aligned features enable more accurate fusion.
Soft Occupancy Confidence Calculation. To compare different methods for computing soft occupancy confidence, we evaluated the performance of tanh and sigmoid functions, each combined with a range-mapping operation to normalize inputs from , into . As illustrated in Table C3, these functions yield comparable results and effectively meet our objectives. However, our exponential-based function achieved a highest mIoU, indicating its superior performance.
| mIoU | IoU | |
| Ours | 18.02 | 44.41 |
| Tanh | 14.85 | 42.31 |
| Sigmoid | 14.99 | 39.57 |
Frame Design and Efficiency. To evaluate our frame design and efficiency, especially in the final feature fusion, we introduce a comparison called ‘Three Fus’, which fuses lifted image, segmentation, and depth features via dual interaction. Table C4 reports the inference time of our voxel fusion module alongside other voxel-based learning modules. Our depth-aware fusion attains a higher mIoU than ThreeFus. Our voxel fusion processes dense voxels in 177.07 ms, comparable to the sparse-voxel modules used in VoxFormer [27] and SGN [38].
| Method | mIoU |
other |
barrier |
bicycle |
bus |
car |
cons. veh. |
motorcycle |
pedestrian |
traffic cone |
trailer |
truck |
drive. surf. |
other flat |
sidewalk |
terrain |
manmade |
vegetation |
| MonoScene [5] | 6.06 | 1.75 | 7.23 | 4.26 | 4.26 | 4.93 | 9.38 | 3.98 | 3.98 | 3.90 | 4.45 | 7.17 | 14.91 | 6.32 | 7.92 | 7.43 | 1.01 | 7.43 |
| OccFormer [54] | 21.93 | 5.95 | 21.93 | 5.85 | 37.83 | 17.87 | 40.44 | 42.43 | 7.36 | 23.88 | 21.81 | 20.98 | 22.38 | 30.70 | 55.35 | 28.36 | 30.70 | 18.0 |
| BEVFormer [30] | 26.88 | 5.94 | 30.29 | 12.32 | 34.40 | 39.17 | 14.44 | 16.45 | 17.22 | 9.27 | 13.90 | 26.36 | 50.99 | 30.96 | 34.66 | 22.73 | 6.76 | 6.97 |
| TPVFormer [16] | 27.83 | 7.22 | 38.90 | 13.67 | 40.78 | 45.90 | 17.23 | 19.99 | 18.85 | 14.30 | 26.69 | 34.17 | 55.65 | 35.47 | 37.55 | 30.70 | 19.40 | 16.78 |
| CTF-Occ [45] | 28.53 | 8.09 | 39.33 | 20.56 | 38.29 | 42.24 | 16.93 | 24.52 | 22.72 | 21.05 | 22.98 | 31.11 | 53.33 | 33.84 | 37.98 | 33.23 | 20.79 | 18.00 |
| RenderOcc [39] | 26.11 | 4.84 | 31.72 | 10.72 | 27.67 | 26.45 | 13.87 | 18.2 | 17.67 | 17.84 | 21.19 | 23.25 | 63.20 | 36.42 | 46.21 | 44.26 | 19.58 | 20.72 |
| Ours | 28.40 | 9.05 | 32.07 | 15.22 | 26.51 | 39.66 | 13.65 | 15.35 | 17.62 | 18.40 | 17.64 | 22.53 | 73.73 | 34.67 | 42.75 | 45.68 | 30.53 | 27.80 |
Fusion Mechanism. To evaluate the design of the voxel fusion module, we compare it with five alternatives: using Conv3D layers (Conv.), using Swin Transformer (Swin.), using as the query with DA (Single DA =img.), using as the query with DA (Single DA. =sem.), and using dual DA with 2 layers (DA 2 layers). As depicted in Table C6, Conv. and Swin. encounter challenges in capturing long-range dependencies, decreasing mIoU by 1.27 and 1.49, respectively. Removing either interaction in baseline also results in mIoU drop. The 2-layer version does not yield any prediction increase but increases inference time to 283.76. In a nutshell, our voxel fusion module strikes a balance between precision and efficiency.
| Inf. Time (ms) | mIoU | IoU | |
| Conv. | 0.16 | 16.75 (-1.27) | 40.55 |
| Swin. | 0.67 | 16.53 (-1.49) | 41.30 |
| Single DA (=Img.) | 134.52 | 16.94 (-1.08) | 44.69 |
| Single DA (=Sem.) | 134.12 | 17.88 (-0.14) | 42.91 |
| 2 Layers DA | 283.76 | 17.48 (-0.54) | 45.16 |
| Ours | 177.07 | 18.02 | 44.41 |
@@ C.2 Supplementary Qualitative Results
To demonstrate the robustness of DSOcc, we present additional visualization results in Fig. C2. DSOcc effectively identifies each vehicle in front of the ego vehicle. It can also detect small objects, such as pedestrians, cyclists, and poles. Furthermore, DSOcc provides robust scene completion when navigating intersections.
C.3 Supplementary Comparison
We also present a quantitative comparison on the Occ3D-nuScenes [45] validation set in Table C5. The dataset comprises 700 training scenes and 150 validation scenes, each annotated within an volume at a voxel resolution for 18 semantic classes. Because Occ3D-nuScenes targets multi-view inputs, unlike our multi-frame setup, we adapt it by disabling multi-segmentation fusion and using only a single frame during DSOcc training. We employ SurroundDepth [48] for depth estimation and train DSOcc on 8 NVIDIA V100 GPUs. Despite these modifications, DSOcc remains competitive, achieving a mIoU of 20.00
@@ Appendix D Limitation
As shown in Fig. C3, when the vehicle rounds a corner, our method sometimes fails to reconstruct the side view because of missing camera features and abrupt changes in vehicle calibration. Our method recognizes this region as the flat other-ground class. This issue could be mitigated by incorporating additional camera viewpoints or by integrating other modalities, such as LiDAR.